How would you begin to solve the equation 3x+x(x-1)=0. I have worked and worked on this problem. I have an answer but when put back in to equation to check. It does not work. My answer is 1 + square root of 11 / 2 , 1 - square root of 11/2.
Answer by ankor@dixie-net.com(22740)

How would you begin to solve the equation 3x+x(x-1)=0
Perform the indicated operations:
3x + x^2 - x = 0
:
x^2 + 3x - x = 0
:
x^2 + 2x = 0
:
Factor out x:
x(x + 2) = 0
:
Giving you two solutions
x = 0
and
x = -2
:
Substituting either solution, confirms the equation